OpenStudy (anonymous):

I need help figuring out how to do this! Help is welcome :) 11. If and object is dropped from a height of 144 ft, the function h(t)=-16t^2+144 gives the height of the object after t seconds. When will the object hit the ground? (1 point) a. 1.5 s b. 3 s c. 6 s d. 9 s

OpenStudy (rulnick):

set h=0

OpenStudy (sidsiddhartha):

h(t)=-16t^2+144 h(t)=0 then 16t^2=144 t=3

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (rulnick):

... so 16t^2=144